# This version: August 29, 2025 )

### v3

@@ 2 Model
We consider a continuum population of students and a set of $J$ schools, $\mathcal{J}:=\{1,\ldots,J\}$, that have capacities $\{q_{1},\ldots,q_{J}\}$ defined in terms of shares of the student population (Azevedo and Leshno, 2016). Denote by $\Omega$ the set of all students in the universe of interest and use $\omega$ to index an individual student type. The student type consists of three objects. First, $Q(\omega)$ denotes the true (strict) preference relation of student $\omega$ over the set of options $\mathcal{J}^{0}:=\mathcal{J}\cup\{0\}$, which includes schools $\mathcal{J}$ and an outside option 0. For example, if $J=2$ and $Q(\omega)=\{1,2,0\}$, then 1 is preferred to 2 (i.e., $1Q(\omega)2$), 1 is preferred to 0 (i.e., $1Q(\omega)0$), and 2 is preferred to 0 (i.e., $2Q(\omega)0$). Let $\mathcal{Q}$ be the set of all strict preference relations over $\mathcal{J}^{0}$ that admit at least one school that is acceptable. A school $j\in\mathcal{J}$ is “acceptable” for student $\omega$ if it is preferred to that student’s outside option, i.e., $jQ(\omega)0$. We define $\bar{Q}$ as the weak preference relation induced by $Q$, i.e., $j\bar{Q}k\Leftrightarrow jQk$ or $j=k$. The second object of the student type is a vector of scores $\bm{\mathrm{R}}(\omega):=(R_{1}(\omega),\ldots,R_{J}(\omega))\in\mathcal{R}\subseteq\mathbb{R}^{J}$, where each school $j$ utilizes $R_{j}$ to rank students for admission. The third and last object, $Y(\omega,d)$, is the potential outcome of student $\omega$ if the student is assigned to option $d\in\mathcal{J}^{0}$. Each student has a potential outcome function $Y(\omega,\cdot)$ that maps from $\mathcal{J}^{0}$ to $\mathcal{Y}\subseteq\mathbb{R}$. We call $\bm{\mathrm{\Gamma}}$ the set of all possible potential outcome functions. The set of all student types is $\Omega:=\mathcal{Q}\times\mathcal{R}\times\bm{\mathrm{\Gamma}}$. In a continuum economy, there is a probability measure $\mathbb{P}$ over $\Omega$ and the Borel $\sigma$-algebra of the product space $\Omega$. We suppress the argument $\omega$ whenever it is unnecessary for ease of notation, e.g., $Y(d)$ vs. $Y(\omega,d)$ and $Q$ vs. $Q(\omega)$.
@@
A “matching” is described by a measurable function $\mu:\Omega\to\mathcal{J}^{0}$ that satisfies two conditions: for every $j\in\mathcal{J}$, (i) the mass of students matched to $j$ is less than or equal to the capacity of school $j$, i.e., $\mathbb{P}\{\omega:\mu(\omega)=j\}\leq q_{j}$; and (ii) the set of students who weakly prefer option $j\in\mathcal{J}^{0}$ over their matching, i.e., $\{\omega:j\bar{Q}(\omega)\mu(\omega)\}$, is an open set.444 Azevedo and Leshno (2016) impose the same condition to rule out multiplicity of stable matchings that differ in a set of types with measure zero. For every student type $\omega$, $\mu(\omega)$ is either the school $j$ to which the student is matched or zero. When $\mu(\omega)=0$, the student is unmatched and takes an outside option. An important definition for this paper is that of stability.
@@
A mechanism $\varphi$ matches students to schools by mapping the students’ scores and submitted preference lists to schools. Student $\omega$ submits a preference list $P(\omega)\subseteq\mathcal{J}$, which is an ordered list of her acceptable schools. For example, for $J=3$, if $Q(\omega)=\{1,2,0,3\}$, then $P(\omega)=\{1,2\}$, as long as the student submits her true list of acceptable schools. The number of schools in $P$, denoted $|P|$, is at least one because everyone participating in the match has at least one acceptable school. As with $\bar{Q}$, we also define $\bar{P}$ as the weak preference relation induced by $P$. A mechanism takes as inputs everyone’s submitted preferences (i.e., a correspondence $P:\Omega\rightrightarrows\mathcal{J}$) and everyone’s scores (i.e., a vector-valued function $\bm{\mathrm{R}}:\Omega\to\mathcal{R}$) and gives rise to a matching function. Formally, $\varphi(P,\bm{\mathrm{R}}):\Omega\to\mathcal{J}^{0}$. We say a mechanism $\varphi$ is strategy proof if, for every student, submitting the true ranking of acceptable schools is a weakly dominant strategy—in other words, any student $\omega$’s misreporting of $P$ never leads to a better option and sometimes leads to a worse option, depending on what other students submit. We say a student is a truth-teller if her $P$ equals her true ranking of acceptable schools. Otherwise, we say she is strategic or not a truth-teller.
@@
This paper considers mechanisms that produce matching functions with a cutoff characterization according to Definition 2. Placement scores $\bm{\mathrm{S}}$ may or may not equal school priority scores $\bm{\mathrm{R}}$. The definition gives researchers the freedom to construct special placement scores $\bm{\mathrm{S}}$ if the mechanism that they consider does not admit cutoff characterization by means of priority scores $\bm{\mathrm{R}}$. The idea behind this definition stems from the logic of the general class of mechanisms of Agarwal and Somaini (2018). Moreover, we assume throughout the paper that, for any $j\in\mathcal{J}$, $S_{j}$ has a distribution that is absolutely continuous with respect to (wrt) the Lebesgue measure and support $\mathcal{S}_{j}$ that contains a closed interval around $c_{j}$. The support of $\bm{\mathrm{S}}$ is $\mathcal{S}$. For any two scores $S_{j}$ and $S_{l}$, we assume that either $\mathbb{P}[S_{j}=S_{l}]=1$ or $\mathbb{P}[f(S_{j})=S_{l}]<1$ for any measurable function $f$. This says that the only deterministic function relating any two scores may be the identity function. Azevedo and Leshno (2016) demonstrate that, if the matching function is stable, the matching has cutoff characterization with $\bm{\mathrm{S}}=\bm{\mathrm{R}}$ and admission cutoffs constructed as follows. For each $j\in\mathcal{J}$, $c_{j}:=\inf\{S_{j}(\omega):\text{ for }\omega\text{ with }\mu(\omega)=j\}$ if some individuals are matched to $j$ or $c_{j}:=\inf\mathcal{S}_{j}$ if nobody is matched to $j$. Many mechanisms produce stable matchings. For example, SD and DA are strategy proof and lead to stable matchings if agents are truth-tellers. Regarding settings where agents are not truth-tellers, e.g., because they face constraints in the submission of $P$, Haeringer and Klijn (2009) study how the SD, DA, Boston, and TTC mechanisms lead to stable matchings.
@@
Having laid out the physical environment of our matching economy, we now motivate the type of causal parameter that concerns this paper. Causal parameters are often used as policy tools for their internal validity—evaluating the impact of historical interventions on outcomes—or for their external validity—assessing the impact of new interventions based on knowledge of historical interventions, that is, constructing counterfactual analyses; see Heckman (2005) for a detailed discussion.
In this paper, we aim to provide strategies for identifying a set of parameters that are not only internally valid but also possess substantial external validity, thereby enabling the evaluation of counterfactual policies. We seek to identify conditional moments of treatment effects $Y(d^{\prime})-Y(d)$ when the econometrician has access to an infinite amount of data and observes the joint distribution of the following random objects: $P(\omega)$, $\bm{\mathrm{S}}(\omega)$, $\mu(\omega)$, and $Y(\omega):=Y(\omega,\mu(\omega))$.555 We abuse the notation and employ the letter $Y$ for both the observed outcome, $Y(\omega)$, and the potential outcome of being assigned to school $d$, $Y(\omega,d)$.
Section A.1 in the appendix provides a detailed discussion of a class of counterfactual analyses that is advocated by Artemov et al. (2023). Analyzing the welfare effects of a policy change in that class relies on identification of the average structural functions (ASF)
@@ 3 Identification with Truthful Reports
In this section, we consider identification of causal effects when all students are truth-tellers, that is, when they submit their true ranking of acceptable schools. We start with truth-telling in order to introduce our approach and notation in a simple behavioral setting before moving to strategic reports in Section 4. Truth-telling is rational in strategy-proof mechanisms. For instance, Dubins and Freedman (1981) and Roth (1982) show that the DA mechanism without constraints on preference submission is strategy proof in the finite economy; Abdulkadiroglu et al. (2015) show the same for the continuum economy.666 See also the work by Azevedo and Budish (2019), who advocate a robust notion of strategyproofness in a large economy. The SD is also strategy proof, as it is a special case of DA. On the other hand, experimental evidence show that some individuals deviate from the rational behavior of truth-telling (Chen and Sönmez (2006), Pais and Pintér (2008)). Thus, we explicitly assume everyone is a truth-teller.
@@
It is straightforward to obtain counterfactual sets of available schools in the case where all schools rely on the same placement score, that is, $S_{j}=S_{1}$ for every $j$. For example, this is the case under the SD mechanism. In this case, the set of feasible schools is all schools with a cutoff below or equal to score $S_{1}$. Note that everyone just above (or just below) cutoff $c_{j}$ has exactly the same set of feasible schools. For someone with $S_{1}\geq c_{j}$, the counterfactual scenario has the score crossing to the left of cutoff $c_{j}$, and school $j$ is dropped from the set of feasible schools. In turn, for someone with $S_{1}<c_{j}$, the counterfactual scenario adds school $j$ to the set of feasible schools. Unlike under SD, agents near and on the same side of a cutoff in DA differ in their sets of feasible schools. It is not immediately obvious which schools appear in their counterfactual sets. In DA, schools use different scores, and these scores may be functions (e.g., weighted averages) of a small set of primitive scores. That is the case in our application with the Chilean data. This makes the joint support of the distribution of scores highly dependent and complicates the counterfactual analysis. Dealing with this complexity is empirically relevant since many real-world higher education assignment mechanisms use DA. See Table 1 Panel B in Fack et al. (2019) for a list of examples.777 A common practice in applied work consists of “cleaning” irrelevant schools from the submitted preference lists in cases where $S_{j}=S_{1}$ for every $j$. For instance, say an individual submits $P=\{1,2,3\}$ and $c_{2}>S_{1}>c_{1}>c_{3}$. Given the cutoff characterization and truth-telling, the matching assignment of this individual is school $1$; the counterfactual assignment when $c_{2}>c_{1}>S_{1}>c_{3}$ is school $3$ even though $2P3$; this is the case because school $2$ has a cutoff higher than the cutoff of school $1$. In this case, the irrelevant school to be cleaned from $P$ is school $2$. The general idea is to remove all schools ranked below $1$ that have cutoffs higher than $c_{1}$. See the description of this practice by Estrada and Gignoux (2017). The practice cannot be used to identify counterfactual assignments in cases where different schools use different placement scores, as under, e.g., the DA mechanism.
@@
In SD, $S_{j}=S_{1}$ for every $j$, and the definition of the budget set above equals $B(\bm{\mathrm{S}})=\{m:S_{1}\geq c_{m}\}$. Suppose we have four schools with cutoffs $c_{1}<c_{2}<c_{3}<c_{4}$. Individuals in this economy have five possible budget sets: $\{0\}$, $\{0,1\}$, $\{0,1,2\}$, $\{0,1,2,3\}$, and $\{0,1,2,3,4\}$. For individuals near cutoff $c_{4}$, the counterfactual budget sets are $B^{-}_{4}(\bm{\mathrm{S}})=\{0,1,2,3\}$ and $B^{+}_{4}(\bm{\mathrm{S}})=\{0,1,2,3,4\}$.888 In SD or in DA with independent placement scores, the definitions of the counterfactual equal $B^{+}_{j}(\bm{\mathrm{S}})=B(\bm{\mathrm{S}})\cup\{m:c_{m}=c_{j}\}$ and $B^{-}_{j}(\bm{\mathrm{S}})=B(\bm{\mathrm{S}})\setminus\{m:c_{m}=c_{j}\}$; moreover, if the cutoffs are unique, $B^{+}_{j}(\bm{\mathrm{S}})=B(\bm{\mathrm{S}})\cup\{j\}$ and $B^{-}_{j}(\bm{\mathrm{S}})=B(\bm{\mathrm{S}})\setminus\{j\}$.
Next, we follow the intuition of Kirkeboen et al. (2016) and define the concept of local preferences, that is, the first- and second-best choices for a marginal individual at any given cutoff. This will later become the control variable in our RD identification strategy with truthful agents.
@@
We say $(j,k)\in\mathcal{J}\times\mathcal{J}$, $j\neq k$, is a comparable pair of alternatives if (i) $c_{j}$ is an interior point of the support $\mathcal{S}_{j}$ and (ii) $\mathbb{P}[Q_{j}=(j,k)|S_{j}=s]$ is bounded away from zero for $s$ in an open neighborhood of $c_{j}$. Finally, we define $\mathcal{P}\subseteq\mathcal{J}\times\mathcal{J}$ as the set of all comparable pairs.
We adopt the convention that comparable pairs do not involve the outside option 0. We do this because it may be hard to interpret the treatment effects of a change from the outside option to a school when the outside option varies across individuals. We do not consider pairs with $j=k$ because the initial school assignment does not change for these individuals. We also exclude pairs $Q_{j}=(j^{\prime},k)$ with $j^{\prime}\neq j$ from $\mathcal{P}$ to avoid redundancy. We may find individuals with $Q_{j}=(j^{\prime},k)$ whenever schools $j$ and $j^{\prime}$ use the same score and have the same cutoff. As the score $S_{j}=S_{j^{\prime}}$ crosses the cutoff $c_{j}=c_{j^{\prime}}$, access is granted to both schools $j$ and $j^{\prime}$, and individuals may differ in their preferences for these schools. Individuals who prefer $j^{\prime}$ will not appear in $\mathcal{P}$ as having $Q_{j}=(j^{\prime},k)$, but they may appear in $\mathcal{P}$ with $Q_{j^{\prime}}=(j^{\prime},k)$.

### v4

@@ 2 Model
We consider a continuum population of students and a set of $J$ schools, $\mathcal{J}:=\{1,\ldots,J\}$, that have capacities $\{q_{1},\ldots,q_{J}\}$ defined in terms of shares of the student population (azevedo2016). Denote by $\Omega$ the set of all students in the universe of interest and use $\omega$ to index an individual student type. The student type consists of three objects. First, $Q(\omega)$ denotes the true (strict) preference relation of student $\omega$ over the set of options $\mathcal{J}^{0}:=\mathcal{J}\cup\{0\}$, which includes schools $\mathcal{J}$ and an outside option 0. For example, if $J=2$ and $Q(\omega)=\{1,2,0\}$, then 1 is preferred to 2 (i.e., $1Q(\omega)2$), 1 is preferred to 0 (i.e., $1Q(\omega)0$), and 2 is preferred to 0 (i.e., $2Q(\omega)0$). Let $\mathcal{Q}$ be the set of all strict preference relations over $\mathcal{J}^{0}$ that admit at least one school that is acceptable. A school $j\in\mathcal{J}$ is “acceptable” for student $\omega$ if it is preferred to that student’s outside option, i.e., $jQ(\omega)0$. We define $\bar{Q}$ as the weak preference relation induced by $Q$, i.e., $j\bar{Q}k\Leftrightarrow jQk$ or $j=k$. The second object of the student type is a vector of scores $\bm{\mathrm{R}}(\omega):=(R_{1}(\omega),\ldots,R_{J}(\omega))\in\mathcal{R}\subseteq\mathbb{R}^{J}$, where each school $j$ utilizes $R_{j}$ to rank students for admission. The third and last object, $Y(\omega,d)$, is the potential outcome of student $\omega$ if the student is assigned to option $d\in\mathcal{J}^{0}$. Each student has a potential outcome function $Y(\omega,\cdot)$ that maps from $\mathcal{J}^{0}$ to $\mathcal{Y}\subseteq\mathbb{R}$. We call $\bm{\mathrm{\Gamma}}$ the set of all possible potential outcome functions. The set of all student types is $\Omega:=\mathcal{Q}\times\mathcal{R}\times\bm{\mathrm{\Gamma}}$. In a continuum economy, there is a probability measure $\mathbb{P}$ over $\Omega$ and the Borel $\sigma$-algebra of the product space $\Omega$. We suppress the argument $\omega$ whenever it is unnecessary for ease of notation, e.g., $Y(d)$ vs. $Y(\omega,d)$ and $Q$ vs. $Q(\omega)$.
@@
A “matching” is described by a measurable function $\mu:\Omega\to\mathcal{J}^{0}$ that satisfies two conditions: for every $j\in\mathcal{J}$, (i) the mass of students matched to $j$ is less than or equal to the capacity of school $j$, i.e., $\mathbb{P}\{\omega:\mu(\omega)=j\}\leq q_{j}$; and (ii) the set of students who weakly prefer option $j\in\mathcal{J}^{0}$ over their matching, i.e., $\{\omega:j\bar{Q}(\omega)\mu(\omega)\}$, is an open set.666azevedo2016 impose the same condition to rule out multiplicity of stable matchings that differ in a set of types with measure zero. For every student type $\omega$, $\mu(\omega)$ is either the school $j$ to which the student is matched or zero. When $\mu(\omega)=0$, the student is unmatched and takes an outside option. An important definition for this paper is that of stability.
@@
A mechanism $\varphi$ matches students to schools by mapping the students’ scores and submitted preference lists to schools. Student $\omega$ submits a preference list $P(\omega)\subseteq\mathcal{J}$, which is an ordered list of her acceptable schools. For example, for $J=3$, if $Q(\omega)=\{1,2,0,3\}$, then $P(\omega)=\{1,2\}$, as long as the student submits her true list of acceptable schools. The number of schools in $P$, denoted $|P|$, is at least one because everyone participating in the match has at least one acceptable school. As with $\bar{Q}$, we also define $\bar{P}$ as the weak preference relation induced by $P$. A mechanism takes as inputs everyone’s submitted preferences (i.e., $P$ which is a mapping from $\Omega$ to an ordered subset of $\mathcal{J}$) and everyone’s scores (i.e., a vector-valued function $\bm{\mathrm{R}}:\Omega\to\mathcal{R}$) and gives rise to a matching function. Formally, $\varphi(P,\bm{\mathrm{R}}):\Omega\to\mathcal{J}^{0}$. We say a mechanism $\varphi$ is strategy proof if, for every student, submitting the true ranking of acceptable schools is a weakly dominant strategy—in other words, any student $\omega$’s misreporting of $P$ never leads to a better option and sometimes leads to a worse option, depending on what other students submit. We say a student is a truth-teller if her $P$ equals her true ranking of acceptable schools. Otherwise, we say she is strategic or not a truth-teller.
@@
This paper considers mechanisms that produce matching functions with a cutoff characterization according to Definition 2. Placement scores $\bm{\mathrm{S}}$ may or may not equal school priority scores $\bm{\mathrm{R}}$. The definition gives researchers the freedom to construct special placement scores $\bm{\mathrm{S}}$ if the mechanism that they consider does not admit cutoff characterization by means of priority scores $\bm{\mathrm{R}}$. The idea behind this definition stems from the logic of the general class of mechanisms of agarwal2018. Moreover, we assume throughout the paper that, for any $j\in\mathcal{J}$, $S_{j}$ has a distribution that is absolutely continuous with respect to (wrt) the Lebesgue measure and support $\mathcal{S}_{j}$ that contains a closed interval around $c_{j}$. The support of $\bm{\mathrm{S}}$ is $\mathcal{S}$. For any two scores $S_{j}$ and $S_{l}$, we assume that either $\mathbb{P}[S_{j}=S_{l}]=1$ or $\mathbb{P}[f(S_{j})=S_{l}]<1$ for any measurable function $f$. This says that the only deterministic function relating any two scores may be the identity function.
azevedo2016 demonstrate that, if the matching function is stable, the matching has cutoff characterization with $\bm{\mathrm{S}}=\bm{\mathrm{R}}$ and admission cutoffs constructed as follows. For each $j\in\mathcal{J}$, $c_{j}:=\inf\{S_{j}(\omega):\text{ for }\omega\text{ with }\mu(\omega)=j\}$ if some individuals are matched to $j$ or $c_{j}:=\inf\mathcal{S}_{j}$ if nobody is matched to $j$. In fact, most empirical work applying RD to matching data in the last 15 years rely on this stability-based method for constructing cutoffs. Many mechanisms produce stable matchings. For example, SD and DA are strategy proof and lead to stable matchings if agents are truth-tellers.
Regarding settings where agents are not truth-tellers, e.g., because they face constraints in the submission of $P$, haeringer2009 and fack2019 show how stability arises in equilibrium with strategic agents. In particular, fack2019 formally test for stability in their empirical context of Parisian schools and do not find evidence against it. One possible threat to stability is agents that make mistakes during submission of their preferences. artemovCheHe2017; Artemov2023 explicitly study stable matching with mistaken agents. They argue that most of the mistakes observed in equilibrium are likely to be payoff-irrelevant in large markets, and thus not hurting stability of the match. Despite the strong support for stability in the literature, we emphasize that Definition 2 allows for the general cutoff characterization of agarwal2018 that does not require stability.
@@
Having laid out the physical environment of our matching economy, we now motivate the type of causal parameter that concerns this paper. Causal parameters are often used as policy tools for their internal validity—evaluating the impact of historical interventions on outcomes—or for their external validity—assessing the impact of new interventions based on knowledge of historical interventions, that is, constructing counterfactual analyses; see heckman2005scientific for a detailed discussion.
In this paper, we aim to provide strategies for identifying a set of parameters that are not only internally valid but also possess substantial external validity, thereby enabling the evaluation of counterfactual policies. We seek to identify conditional moments of treatment effects $Y(d^{\prime})-Y(d)$ when the econometrician has access to an infinite amount of data and observes the joint distribution of the following random objects: $P(\omega)$, $\bm{\mathrm{S}}(\omega)$, $\mu(\omega)$, and $Y(\omega):=Y(\omega,\mu(\omega))$. [^7] We abuse the notation and employ the letter $Y$ for both the observed outcome, $Y(\omega)$, and the potential outcome of being assigned to school $d$, $Y(\omega,d)$.
Section A.1 in the appendix provides a detailed discussion of a class of counterfactual analyses that is advocated by Artemov2023. Analyzing the welfare effects of a policy change in that class relies on identification of the average structural functions (ASF)
@@ 3 Identification with Truthful Reports
In this section, we consider identification of causal effects when all students are truth-tellers, that is, when they submit their true ranking of acceptable schools. We start with truth-telling in order to introduce our approach and notation in a simple behavioral setting before moving to strategic reports in Section 4.
@@
It is straightforward to obtain counterfactual sets of available schools in the case where all schools rely on the same placement score, that is, $S_{j}=S_{1}$ for every $j$. For example, this is the case under the SD mechanism. In this case, the set of feasible schools is all schools with a cutoff below or equal to score $S_{1}$. Note that everyone just above (or just below) cutoff $c_{j}$ has exactly the same set of feasible schools. For someone with $S_{1}\geq c_{j}$, the counterfactual scenario has the score crossing to the left of cutoff $c_{j}$, and school $j$ is dropped from the set of feasible schools. In turn, for someone with $S_{1}<c_{j}$, the counterfactual scenario adds school $j$ to the set of feasible schools. Unlike under SD, agents near and on the same side of a cutoff in DA differ in their sets of feasible schools. It is not immediately obvious which schools appear in their counterfactual sets. In DA, schools use different scores, and these scores may be functions (e.g., weighted averages) of a small set of primitive scores. That is the case in our application with the Chilean data. This makes the joint support of the distribution of scores highly dependent and complicates the counterfactual analysis. Dealing with this complexity is empirically relevant since many real-world higher education assignment mechanisms use DA. See Table 1 Panel B in fack2019 for a list of examples. [^8] A common practice in applied work consists of “cleaning” irrelevant schools from the submitted preference lists in cases where $S_{j}=S_{1}$ for every $j$. For instance, say an individual submits $P=\{1,2,3\}$ and $c_{2}>S_{1}>c_{1}>c_{3}$. Given the cutoff characterization and truth-telling, the matching assignment of this individual is school $1$; the counterfactual assignment when $c_{2}>c_{1}>S_{1}>c_{3}$ is school $3$ even though $2P3$; this is the case because school $2$ has a cutoff higher than the cutoff of school $1$. In this case, the irrelevant school to be cleaned from $P$ is school $2$. The general idea is to remove all schools ranked below $1$ that have cutoffs higher than $c_{1}$. See the description of this practice by estrada2017. The practice cannot be used to identify counterfactual assignments in cases where different schools use different placement scores, as under, e.g., the DA mechanism.
@@
In SD, $S_{j}=S_{1}$ for every $j$, and the definition of the budget set above equals $B(\bm{\mathrm{S}})=\{0\}\cup\{m:S_{1}\geq c_{m}\}$. Suppose we have four schools with cutoffs $c_{1}<c_{2}<c_{3}<c_{4}$. Individuals in this economy have five possible budget sets: $\{0\}$, $\{0,1\}$, $\{0,1,2\}$, $\{0,1,2,3\}$, and $\{0,1,2,3,4\}$. For individuals near cutoff $c_{4}$, the counterfactual budget sets are $B^{-}_{4}(\bm{\mathrm{S}})=\{0,1,2,3\}$ and $B^{+}_{4}(\bm{\mathrm{S}})=\{0,1,2,3,4\}$. [^9] In SD or in DA with independent placement scores, the definitions of the counterfactual equal $B^{+}_{j}(\bm{\mathrm{S}})=B(\bm{\mathrm{S}})\cup\{m:c_{m}=c_{j}\}$ and $B^{-}_{j}(\bm{\mathrm{S}})=B(\bm{\mathrm{S}})\setminus\{m:c_{m}=c_{j}\}$; moreover, if the cutoffs are unique, $B^{+}_{j}(\bm{\mathrm{S}})=B(\bm{\mathrm{S}})\cup\{j\}$ and $B^{-}_{j}(\bm{\mathrm{S}})=B(\bm{\mathrm{S}})\setminus\{j\}$.
Next, we follow the intuition of kirkeboen2016 and define the concept of local preferences, that is, the first- and second-best choices for a marginal individual at any given cutoff. This will later become the control variable in our RD identification strategy with truthful agents.
@@
We say $(j,k)\in\mathcal{J}\times\mathcal{J}$, $j\neq k$, is a comparable pair of alternatives if (i) $c_{j}$ is an interior point of the support $\mathcal{S}_{j}$ and (ii) $\mathbb{P}[Q_{j}=(j,k)|S_{j}=s]$ is bounded away from zero for $s$ in an open neighborhood of $c_{j}$. Finally, we define $\mathcal{P}\subseteq\mathcal{J}\times\mathcal{J}$ as the set of all comparable pairs. [^10] We adopt the convention that comparable pairs do not involve the outside option 0. We do this because it may be hard to interpret the treatment effects of a change from the outside option to a school when the outside option varies across individuals. We do not consider pairs with $j=k$ because the initial school assignment does not change for these individuals. We also exclude pairs $Q_{j}=(j^{\prime},k)$ with $j^{\prime}\neq j$ from $\mathcal{P}$ to avoid redundancy. We may find individuals with $Q_{j}=(j^{\prime},k)$ whenever schools $j$ and $j^{\prime}$ use the same score and have the same cutoff. As the score $S_{j}=S_{j^{\prime}}$ crosses the cutoff $c_{j}=c_{j^{\prime}}$, access is granted to both schools $j$ and $j^{\prime}$, and individuals may differ in their preferences for these schools. Individuals who prefer $j^{\prime}$ will not appear in $\mathcal{P}$ as having $Q_{j}=(j^{\prime},k)$, but they may appear in $\mathcal{P}$ with $Q_{j^{\prime}}=(j^{\prime},k)$.